# Evidence of Memory Effects in the Dynamics of Two-Level System Defect Ensembles Using Broadband, Cryogenic Transient Dielectric Spectroscopy

### v1

@@ Abstract
Two-level system defects (TLSs) in dielectrics are believed to be the primary source of decoherence in superconducting circuits. However, the precise origin of these defects, including their atomistic morphology, frequency distributions, and dipole moments, remains relatively unexplored. A significant reason for this lack of understanding is that TLS defects are currently probed with resonators or qubits that require complex multi-stage fabrication processes. These probes are packaged, cooled in a dilution refrigerator, and measured to infer their corresponding coherence times. However, even frequency-tunable probes only gather information about TLS defects within their mode volume and a narrow bandwidth around their resonance frequency. As a result, a detailed understanding of the atomistic origins of TLS defects in isolated materials and interfaces is currently lacking, making it challenging to design the next generation of superconducting devices and dielectric materials from first principles. Here, we introduce a novel approach that enables probing the cryogenic dielectric response of TLS defect ensembles using a broadband 3D waveguide. This technique provides evidence of memory effects in the dynamics of TLS defect ensembles that may enable the extraction of critical information about the origins of the underlying TLS defects and their interactions. Importantly, this novel spectroscopy technique can be used at all stages of fabrication and ultimately inform mitigation strategies to reduce the impact of TLS defects on superconducting circuits. As a result, this technique will help usher in a new understanding of the atomistic origin of TLS defects, which will have profound consequences for quantum information science and the broader materials science community.
@@ I Introduction
Two-level system (TLS) defects are essential in understanding anomalous low-temperature properties of disordered solids, and they play a critical role in limiting the coherence of quantum technologies such as superconducting qubits and resonators. At low temperatures, many amorphous solids exhibit a thermodynamic behavior that deviates significantly from the predictions of Debye theory, which describes the specific heat of crystalline materials as scaling with $T^{3}$. In contrast, experiments in the early 1970s revealed that glassy materials display a nearly linear temperature dependence of the specific heat at cryogenic temperatures [1]. To explain this anomalous behavior, Anderson, Halperin, and Varma [2], and independently Phillips [3], introduced the standard tunneling model (Fig. 1a). This model posits that structural disorder in dielectrics can give rise to atoms or groups of atoms residing in effective double-well potentials, allowing for quantum tunneling between two nearly degenerate configurations.
@@
Figure 1: Overview of TLS defects and how they can be probed. a, Double-well representation of TLS defects described by the standard tunneling model. b, Candidate TLS defects and their potential locations in superconducting circuits. c, Comparison of traditional TLS defect spectroscopy using 2D qubits and resonators and our proposed broadband 3D waveguide approach. Image in a is adapted from [4].
@@
These tunneling systems are ubiquitous and form an ensemble of TLS defects, whose low-energy excitations account for the observed linear temperature dependence of the specific heat capacity. Early models treated TLS defects as non-interacting and uniformly distributed in frequency, which sufficed to explain the specific heat measurements. However, studies have shown that TLS defects can interact, leading to spectral diffusion, clustering, and nonlinear loss [5, 6, 7, 8, 9, 10, 11].
@@
While the notion of a TLS defect was initially developed to explain the thermodynamic properties of glassy materials, such defects have recently regained relevance in solid-state quantum computing technologies. This renewed interest is especially evident in superconducting circuits, which have become a leading platform primarily due to their fast gate times, design flexibility, and compatibility with advanced semiconductor fabrication techniques. However, their decoherence times are relatively short compared to other promising quantum technologies, such as neutral atoms, trapped ions, or spin qubits. From measurements of bulk loss tangents [12, 13], and modeling of the Purcell effect [14, 15, 16], it became clear that other sources of loss must be present. Pioneering studies examining the impact of electric fields and mechanical strain on qubit coherence provided explicit evidence of TLS defects in superconducting circuits [17, 18, 19, 20, 21]. Today, it is widely believed that TLS defects are the dominant source of decoherence of superconducting circuits and that they likely reside at surfaces and interfaces [22, 23, 24, 25, 26] (Fig. 1b) with no clear consensus as to their specific atomistic origin.
TLS defects are typically probed via photonic or acoustic resonators [27, 28, 29, 30, 31] and qubits [12, 32, 17, 18, 19, 20]. These approaches provide high cooperativity but preferentially detect TLS defects with dipole moments aligned with the directly integrated probes, and are restricted to small bandwidths and mode volumes (Fig. 1c). Despite this limitation, several studies have demonstrated the existence of TLS-TLS interactions [33, 34, 19], and non-Markovian dynamics [35, 36], leading to questions about the nature of TLS-TLS interactions in the absence of strongly coupled probes [37].
@@
In addition to their impact on the decoherence of superconducting qubits, TLS defects also influence the fidelity of qubit readout. In particular, TLS defects can manifest in Josephson traveling wave parametric amplifiers (JTWPAs), which are crucial for high-fidelity qubit readout due to their quantum-limited amplification and several GHz of bandwidth [38]. However, recent experiments show that high-power pulses can excite long-lived dielectric echoes in JTWPAs, which are attributed to TLS defect ensembles [39, 40]. These echoes impair the readout fidelity of experiments utilizing JTWPAs (and likely other parametric amplifiers), motivating careful characterization of the transient dynamics of interacting TLS defect ensembles.
@@
Here, we propose a novel, modular, noninvasive technique that enables broadband dielectric spectroscopy of TLS defect ensembles in isolated materials and material stacks. Through systematic, longitudinal studies, this technique will provide essential data that will shed light on the atomistic nature of TLS defects and their properties [41]. Such insights will inform the design of the next generation of quantum hardware from a first-principles understanding of TLS defects, something which was recently highlighted as a pressing need for the superconducting qubit community by the Google Quantum AI Hardware team [42].
@@ II CRYOGENIC DIELECTRIC SPECTROSCOPY
Figure 2: Broadband waveguide design. a, Photograph of upper and lower WR-229 to SMA adapter components and clamp with sapphire samples. These components are assembled to create the closed 3D aluminum waveguide used in this work. b, HFSS simulated and measured transmission ($S_{21}$) and reflection ($S_{11}$) spectra of the waveguide containing samples at room temperature, demonstrating a waveguide cutoff of 2.8 GHz and broadband transmission from 2.8-6 GHz.
To probe the dielectric response in the microwave frequency domain, we leverage a broadband 3D waveguide (Fig. 2a). In contrast to qubit and resonator-based probes, this method allows coupling to ensembles of TLS defects over a broad frequency range (Fig.1c). As described in Appendix A, the closed waveguide comprises two WR-229 to coaxial cable adapters joined by a sample clamp. This modular approach allows us to easily measure samples of various sizes, shapes, and material compositions (Fig. 2b). Importantly, these features distinguish our approach from existing narrow-band qubit and resonator-based probes, which require samples to go through a full fabrication cycle, followed by wire bonding and packaging, making it challenging to perform additional material processing and characterization.
@@
Applying an external electric field to dielectric materials induces a time-dependent polarization that reflects the collective dielectric response. This response is dominated by a distribution of atomic-scale TLS defects at low temperatures, whose energy splittings and coupling strengths vary widely due to structural disorder. Within the linear response regime, the relationship between the applied field and the induced polarization is captured by the dynamical susceptibility of the material, which encodes both in- and out-of-phase components of the TLS ensemble response. In our experiment, the response can extend beyond the linear region when the system is influenced by strong, broadband, or pulsed electric fields, making the susceptibility dependent on amplitude, leading to higher-order contributions. In this context, we use the dynamical susceptibility framework as a conceptual tool to interpret the response, even though saturation and non-linear effects play an essential role. After the drive is turned off, the system enters a transient regime, where the energy absorbed by TLS defects begins to dissipate. By monitoring the subsequent time evolution of the polarization using homodyne detection, we perform broadband transient dielectric spectroscopy, which reveals signatures of the TLS defects dynamics.
We consider a single TLS defect characterized by an energy detuning $\varepsilon_{j}$ and a tunneling amplitude $\Delta_{j}$, as illustrated in Fig. 1a. The effective Hamiltonian for the $j$-th TLS defect in the position basis $\{|L\rangle,|R\rangle\}$ can be written as [19]
@@
Next, we introduce the interaction of the TLS defect ensemble with an external, time-dependent electric field $\mathbf{E}(t)$. The coupling occurs via the electric dipole moment $\mathbf{p}_{j}$ of the TLS defect. Choosing the dipole moment direction as the quantization axis (here aligned along $\hat{\mathbf{x}}$), the collective polarization operator for an ensemble of $N$ TLS defects can be defined as
@@
In typical experimental setups, the electromagnetic field is polarized and can be represented as
@@
This formulation captures the static energy structure of the TLS defect ensemble and its collective coupling to the external, oscillating electric field. However, in practice, the dynamical response includes additional features such as dynamical Stark shifts and nonlinearities [43]. In previous narrow-band probe experiments using qubits and resonators, the driving is typically focused on the qubit or resonator, and such direct driving of TLS defects is ignored. The driving term in Eq. 6 is absent in the standard tunneling model.
@@
$\chi^{\prime\prime}(\omega)\equiv\operatorname{Im}\chi(\omega)$ (12)
When the electromagnetic field is on, the loss described by Eq. 12 causes a decrease in the measured transmitted signal. However, when this drive is turned off and we enter the transient regime, the energy absorbed from the drive must now be dissipated. Given the insulating nature of the sapphire sample and the superconducting waveguide, a portion of the absorbed energy will be re-emitted and ultimately measured from the output signal of the waveguide. Thus, transient dielectric spectroscopy gives us a direct measure of $\chi^{\prime\prime}(\omega)$ without an input drive. To extract the radiated field to learn about $\chi^{\prime\prime}(\omega)$, we perform homodyne detection, which involves extracting the classical output field quadratures $I(t)$ and $Q(t)$ and computing the amplitude $A(t)=\sqrt{I^{2}(t)+Q^{2}(t)}$. By input-output theory [45], the output field is proportional to the dipole response,
@@
Thus, following Novotny et al. [46], we can express the imaginary part of susceptibility using the experimentally obtained intensity $\mathcal{I}(t)\propto A^{2}(t)$ as
@@
In principle, transient dielectric spectroscopy can be performed on any dielectric sample at any temperature. Furthermore, even metallic or superconducting samples with oxide layers can be probed in reflection. Here, we will be primarily interested in the dielectric response of TLS defects, which solely arise at cryogenic temperatures, as described in Sec. I. Therefore, to extract information about the TLS defects, we perform cryogenic dielectric spectroscopy on our samples by inserting them into the waveguide (Fig. 2a) and cooling it to 10 mK using a dilution refrigerator. As electromagnetic waves propagate through the waveguide and interact with the samples, they polarize the electric dipoles of the TLS defects, revealing their intrinsic dielectric response. This work considers cryogenic dielectric spectroscopy in both driven and transient regimes: driven dielectric spectroscopy monitors the response during the pulse, while transient dielectric spectroscopy monitors the post-pulse region (also known as the ring-down).
@@
To probe a driven dielectric response, we generate 30-ns square pulses in the 3-5 GHz band, similar to those employed for superconducting-qubit gate operations. The simple square pulses we explore here are synthesized directly on an FPGA (Field Programmable Gate Array) board (AMD RFSoC 4x2) using the Quantum Instrumentation Control Kit (QICK) software package [47]. We send the pulses through a room-temperature amplifier (Minicircuits CMA-83LN+) and into a dilution refrigerator (Bluefors LD400) where they are attenuated and ultimately reach the 3D waveguide, exciting the sample (See setup diagram in Appendix E). The resultant signal is amplified using cryogenic high-electron-mobility transistor amplifiers from Low Noise Factory before being measured on the RFSoC.
@@
Figure 3: Dielectric spectroscopy of different samples. We perform dielectric spectroscopy using a 30 ns pulse (marked by vertical black dashed lines) followed by a long readout with a total measurement time of 0.8 $\mu$s. a, Control experiment at room temperature with solvent-cleaned sapphire samples, showing no detectable transient dielectric response. The horizontal band around 4.4 GHz is due to readout electronics, which exhibit a slightly higher noise floor in that frequency range. b, Same setup as in a, measured at 10 mK (as with c-d), revealing coherent transient dielectric response features arising from TLS defects. c, The same samples as b with a 2 nm aluminum oxide layer deposited via atomic layer deposition (ALD). d, The same samples as b with spin-coated Shipley 1813 photoresist. Transient dielectric response in c-d shows prominent features in the ring-down, which we attribute to increased TLS defect density. Subpanels i show the average driven dielectric response and the extracted lifetimes ($\tau$) of the transient dielectric response, which is fitted from the ring-downs, considering a simple exponential function. Regions with large $\tau$ exhibit dips in the driven dielectric response, as expected. Subpanels ii show the driven and transient dielectric response for various samples with the color scale clipped to reveal weaker transient features. In b,ii, we include a vertical white line to indicate the full width at half-maximum of the pulse, $\Delta$f, whose FFT is shown explicitly in the inset and has approximately the same bandwidth as the prominent spectral features in the transient dielectric response. Note that the electric field at the sample plane is frequency dependent and can be calibrated by reducing signal amplitude at certain frequency ranges, as described in Appendix D. For the results displayed in this work, we apply an even amplitude drive across the span to capture maximum response.
@@
Following the cessation of the external drive, the induced polarization in the medium relaxes, leading to a coherent emission in the transient dielectric response, which imprints the temporal decay of the TLS defect ensemble on the outgoing field. A strong time-dependent drive may also modify the phonon-induced decay rate of a TLS by hybridizing its states (dressed states), thereby shifting transition frequencies and enabling decay through induced sidebands. This may alter the coupling to the spectral density of the phonon bath, leading to either enhanced or suppressed relaxation depending on the drive strength. Although we do not investigate this effect here, the broadband nature of our technique should enable its detailed study in future experiments.

### v2

@@ Abstract
Two-level system (TLS) defects in dielectrics are a major source of decoherence in superconducting circuits, yet their atomistic origin, frequency distribution, and dipole moments remain poorly understood. Current probes, which are predominantly based on qubits or resonators, require complex fabrication and only measure defects within a narrow frequency band and limited mode volume, hindering direct insight into TLS defect behaviour in isolated materials and interfaces. Here, we introduce a broadband 3D waveguide spectroscopy technique that enables cryogenic probing of ensembles of TLS defects that we call Broadband Cryogenic Transient Dielectric Spectroscopy (BCTDS). Complementary to the dielectric dipper method, this approach probes a broader spectrum and reveals interference of drive-induced sidebands of the ensembles of TLS defects. The broadband and power-tunable nature of BCTDS makes it especially well-suited to the study of dressed-state physics in driven ensembles of TLS defects, including multi-photon processes and sideband-resolved dynamics. Additionally, BCTDS enables the identification of eigenmode frequencies of the undriven ensembles of TLS defects through characteristic V-shaped features obtained via Fourier analysis of time-domain signals, and shows evidence of memory effects arising from interactions and the broadband nature of our approach. Crucially, our method is modular and can be applied throughout the device fabrication process, informing mitigation strategies and advancing the design of low-loss materials with broad implications for quantum technologies and materials science.
@@ I Introduction
TLS defects are necessary to describe anomalous low-temperature properties of disordered solids, and they play a critical role in limiting the coherence of quantum technologies such as superconducting qubits and resonators. At low temperatures, many amorphous solids exhibit a thermodynamic behavior that deviates significantly from the predictions of Debye theory, which describes the specific heat of crystalline materials as scaling with $T^{3}$. In contrast, experiments in the early 1970s revealed that glassy materials display a nearly linear temperature dependence of the specific heat at cryogenic temperatures [1]. To describe this low temperature behavior, Anderson, Halperin, and Varma [2], and independently Phillips [3], introduced the standard tunneling model, which treats disordered materials as a collection of effective double-well potentials where quantum tunneling can occur (Fig. 1a).
@@
Figure 1: Overview of TLS defects and how they can be probed. a, Double-well representation of TLS defects under a drive $\Omega$, described by the standard tunneling model. b, Candidate TLS defects and their likely locations in superconducting circuits. c, Comparison of traditional TLS defect spectroscopy using 2D qubits and resonators and our proposed broadband 3D waveguide approach. The concept of image b is inspired by [4].
@@
Today, it is widely believed that individual TLS defects are ubiquitous and form an ensemble whose collective low-energy excitations account for the observed linear temperature dependence of the specific heat capacity. Early models treated TLS defects as non-interacting and uniformly distributed in frequency, which sufficed to explain the specific heat capacity measurements. However, studies have shown that TLS defects can interact, leading to spectral diffusion, clustering, and nonlinear loss [5, 6, 7, 8, 9, 10, 11].
@@
While the notion of a TLS defect was initially developed to explain the thermodynamic properties of glassy materials, such defects have recently regained relevance in solid-state quantum computing technologies. This renewed interest is especially evident in superconducting circuits, which have become a leading platform for quantum computing primarily due to their fast gate times, design flexibility, and compatibility with advanced semiconductor fabrication techniques. However, their decoherence times are relatively short compared to other promising quantum technologies, such as neutral atoms, trapped ions, and spin qubits. From measurements of bulk loss tangents [12, 13], and modeling of the Purcell effect [14, 15, 16], it has become clear that other sources of loss must be present. Furthermore, measurements of qubit coherence while changing electric field and mechanical strain provide explicit evidence of TLS defects in superconducting circuits [17, 18, 19, 4, 20]. Today, it is widely believed that the dominant source of decoherence of superconducting circuits is TLS defects that reside on surfaces and interfaces with no clear consensus as to their specific atomistic origin [21, 22, 23, 24, 25] (Fig. 1b).
TLS defects are typically probed via photonic or acoustic resonators [26, 27, 28, 29, 30] and qubits [12, 31, 17, 18, 19, 4]. These approaches provide high cooperativity but preferentially detect TLS defects with dipole moments aligned with the directly integrated probes, and are restricted to small bandwidths and mode volumes (Fig. 1c). Despite these limitations, several studies have demonstrated the existence of TLS-TLS interactions [32, 33, 19], and non-Markovian dynamics [34, 35], leading to questions about the nature of TLS defect interactions in the absence of strongly coupled probes [36]. In the spirit of the Mollow triplet, our setup enables the observation of dressed-state features in interacting ensembles of TLS defects, where the underlying spectral structure can reflect many-body dynamics rather than single-emitter coherence [37]. Furthermore, our method can be extended to accurately map the frequency and dipole moments of ensembles of TLS defects by employing a polarization-dependent drive. While conventional techniques such as electron spin resonance (ESR) focus on detecting net magnetization changes associated with spin-active defects using narrowband field excitation, BCTDS enables parallel, broadband interrogation of both spin and non-spin defects by directly measuring their transient electromagnetic and dielectric responses.
@@
In addition to their impact on the decoherence of superconducting qubits, TLS defects also influence the fidelity of qubit readout. In particular, TLS defects can manifest in Josephson traveling wave parametric amplifiers (JTWPAs), which are crucial for high-fidelity qubit readout due to their quantum-limited amplification and several GHz of bandwidth [38, 39]. However, recent experiments show that high-power pulses can excite long-lived dielectric echoes in JTWPAs, which are attributed to ensembles of TLS defects [40, 41]. These echoes impair the readout fidelity of experiments utilizing JTWPAs (and likely other parametric amplifiers), motivating careful characterization of the transient dynamics of interacting ensembles of TLS defects.
@@
Here, we propose a modular, non-invasive technique that enables BCTDS of ensembles of TLS defects in isolated materials and devices. Through systematic and longitudinal studies, this technique will provide essential data that will shed light on the atomistic nature of TLS defects and their properties [42]. Such insights will inform the design of the next generation of quantum hardware from a first-principles understanding of TLS defects, something which was recently highlighted as a pressing need for the superconducting qubit community by the Google Quantum AI hardware team [43].
@@ II CRYOGENIC DIELECTRIC SPECTROSCOPY
Figure 2: Broadband waveguide design. a, Photograph of upper and lower WR-229 to SMA adapter components and clamp with sapphire samples (marked by the black box). These components are assembled to create the closed 3D aluminum waveguide used in this work. b, HFSS simulated and measured transmission ($S_{21}$) and reflection ($S_{11}$) spectra of the waveguide containing samples at room temperature, demonstrating a waveguide cutoff of around 3 GHz and broadband transmission from 3-6 GHz.
To probe the dielectric response in the microwave frequency domain, we leverage a broadband 3D waveguide (Fig. 2a). In contrast to qubit and resonator-based probes, this method allows for coupling to ensembles of TLS defects over a broad frequency range (Fig. 1c). As described in Appendix  A, the closed waveguide comprises two WR-229 to coaxial cable adapters joined by a sample clamp. This modular approach allows us to easily measure samples of various sizes, shapes, and material compositions over a broadband frequency range of 3-6 GHz (Fig. 2b). Importantly, these features distinguish our approach from existing narrow-band qubit and resonator-based probes, which require samples to go through a full fabrication cycle, followed by wire bonding and packaging, making it nearly impossible to perform additional material processing and characterization. Furthermore, this constraint makes the development of new fabrication recipes wasteful and inefficient.
@@
Applying an external time-dependent electric field to dielectric materials induces a polarization that reflects the collective dielectric response. This response is dominated by a distribution of atomic-scale TLS defects at low temperatures, whose energy splittings and coupling strengths vary widely due to structural disorder. Within the linear response regime, the relationship between the applied field and the induced polarization is captured by the dynamical susceptibility of the material, which encodes both in- and out-of-phase components of the ensembles of TLS defects’ response. In our experiment, the response can extend beyond the linear regime when the system is influenced by strong, broadband, or pulsed electric fields, making the susceptibility dependent on amplitude, leading to higher-order contributions to the susceptibility. In this context, we use the dynamical susceptibility framework as a conceptual tool to interpret the response, even though saturation and non-linear effects likely play a significant role. After the drive is turned off, the system enters a transient regime, where the energy absorbed by TLS defects begins to dissipate. By monitoring the subsequent time evolution of the polarization-dependent homodyne signal, we perform BCTDS, which reveals signatures of the TLS defects dynamics.
To understand this system, we first consider a single TLS defect characterized by an energy detuning $\varepsilon_{j}$ and a tunneling amplitude $\Delta_{j}$, as illustrated in Fig. 1a. The effective Hamiltonian for the $j$-th TLS defect in the position basis $\{|L\rangle,|R\rangle\}$ can be written as [19]
@@
Next, we introduce the interaction of the ensembles of TLS defects with an external, time-dependent electric field $\mathbf{E}(t)$. The coupling occurs via the electric dipole moment $\mathbf{p}_{j}$ of the TLS defect [37]. Choosing the dipole moment direction as the quantization axis (here aligned along $\hat{\mathbf{x}}$), the collective polarization operator for an ensemble of $N$ TLS defects can be defined as
@@
In our experimental setup, the electromagnetic field is polarized and can be represented as
@@
This formulation captures the static energy structure of the ensembles of TLS defects and its collective coupling to the external, oscillating electric field. However, in practice, the dynamical response includes additional features such as Stark shifts and nonlinearities [44]. Unlike prior narrow-band experiments that probe TLS defects indirectly using a qubit or resonator, our approach directly drives the entire ensemble of TLS defects. This motivates including the drive term in Eq. 6 within the standard tunneling model.
@@
$\chi^{\prime\prime}(\omega)\equiv\operatorname{Im}\chi(\omega).$ (12)
When the electromagnetic field interacts with the insulating sapphire samples, the loss described by Eq. 12 causes a decrease in the measured transmitted signal. However, when this drive is turned off and we enter the transient regime, the energy absorbed from the drive must now be re-emitted and ultimately measured from the output signal of the superconducting waveguide. To extract the radiated field, we perform homodyne detection, which involves extracting the classical output field quadratures $I(t)$ and $Q(t)$ and computing the amplitude $A(t)=\sqrt{I^{2}(t)+Q^{2}(t)}$. By input-output theory [46], the output field must be proportional to the dipole response,
@@
Thus, following Novotny et. al. [47], we can express the imaginary part of susceptibility, $\chi^{\prime\prime}$, using the experimentally obtained intensity $\mathcal{I}(t)\propto A^{2}(t)$ as
@@
In principle, BCTDS can be performed on any dielectric sample at any temperature. Furthermore, even metallic or superconducting samples with oxide layers can be probed in reflection (Appendix  D, Fig. 9). Here, we will be primarily interested in the dielectric response of TLS defects, which solely arise at cryogenic temperatures, as described in Sec. I. Therefore, to extract information about the TLS defects, we perform BCTDS on our samples by inserting them into the waveguide (Fig. 2a) and cooling them to 10 mK using a dilution refrigerator. As mentioned in the previous section, driven dielectric spectroscopy monitors the response during the pulse, while transient dielectric spectroscopy monitors the post-pulse region (also known as the ring-down).
@@
To probe a driven dielectric response, we generate square pulses in the 3-5 GHz band, similar to those employed for superconducting-qubit gate operations. The simple square pulses we explore here are synthesized directly on an FPGA (Field Programmable Gate Array) board (AMD RFSoC 4x2) using the Quantum Instrumentation Control Kit (QICK) software package [48]. We send the pulses through a room-temperature amplifier (Minicircuits CMA-83LN+) and into a dilution refrigerator (Bluefors LD400) where they are attenuated and ultimately reach the 3D waveguide and interact with the sample (See setup diagram in Appendix  F). The resultant signal is amplified using cryogenic high-electron-mobility transistor amplifiers from Low Noise Factory before being measured on the RFSoC.
@@
Figure 3: Dielectric spectroscopy of different samples. We perform dielectric spectroscopy using a 30 ns pulse (marked by vertical black dashed lines) followed by a long readout with a total measurement time of 0.8 $\mu$s. Control experiment at room temperature for empty waveguide (a) and waveguide with solvent-cleaned sapphire samples (b) shows no detectable transient dielectric response. The horizontal band around 4.4 GHz is due to readout electronics, which exhibit a slightly higher noise floor in that frequency range. Measurements c-f are performed at 10 mK, revealing coherent transient dielectric response features arising from TLS defects. c, Same setup as a, with no samples mounted in the waveguide. d, Same setup as b with solvent cleaned sapphire samples. e, The same samples as b with a 2 nm aluminum oxide layer deposited via atomic layer deposition (ALD). f, The same samples as b with spin-coated Shipley 1813 photoresist. Transient dielectric response in e-f shows prominent features in the ring-down, which we attribute to increased TLS defect density. Subpanels i show the average driven dielectric response and the extracted lifetimes ($\tau$) of the transient dielectric response, which is fitted from the ring-downs, considering a simple exponential function. Regions with large $\tau$ exhibit dips in the driven dielectric response, as expected. Subpanels ii show the driven and transient dielectric response for various samples with the color scale clipped to reveal weaker transient features. In ii, we include a vertical white line to indicate the full width at half-maximum of the pulse, $\Delta$f, whose FFT is shown explicitly in the inset in dii, and has approximately the same bandwidth as the prominent spectral features in the transient dielectric response. Note that the electric field at the sample plane is frequency dependent and can be calibrated by reducing signal amplitude at certain frequency ranges, as described in Appendix  E. This calibration is not utilized in this data to demonstrate the maximum response.
@@
Following the cessation of the external drive, the induced polarization in the medium relaxes, leading to a coherent emission in the transient dielectric response, which imprints the temporal decay of the ensembles of TLS defects on the outgoing field, allowing us to perform BCTDS. A strong time-dependent drive may also modify the phonon-induced decay rate of a TLS defect by hybridizing its states (dressed states), thereby shifting transition frequencies and enabling decay through induced sidebands. This may alter the coupling to the spectral density of the phonon bath, leading to either enhanced or suppressed relaxation depending on the drive strength. Although we do not investigate this effect here, the broadband nature of our technique should enable its detailed study in future experiments.